OpenStudy (anonymous):

My grand daughter emailed me this and it has been so long since I have done Algebra that I am in a quandry. two consecutive integers have the property that one integer times four times the other equals 224. Can someone help?

OpenStudy (anonymous):

Let one integer by x then as the other integer is consecutive so it must be x+1,by the problem, \[x \times 4 (x+1) = 224 \Rightarrow 4x^2+4x-224= 4 (x-7) (8+x)\] So,we get x =7,-8,ignoring the negative value we get x=4. Lets check : \[ 7 \times 4 (7+1) = 224 \] Looks like we got the correct answer! :)